This is the question in my textbook:
Fran and Winston have a combined income of $80 000. One quarter of Winston's income is the same as one-sixth of Fran's income. How much does each person earn?

Answer by mananth(16946)

combined income = 80000
Let Fran's income be x
Winston's income will be 80000-x
...
1/4 * (80000-x)= 1/6 * x
20000 -x/4= x/6
add x/4
20000 = x/6+x/4
LCD =12
20000=(2x+3x)/12
multiply by 12
12*20000 = 5x
/5
12*4000 =x
x=48000
Fran's income = $ 48,000
Winston's income - $32,000